OpenStudy (anonymous):

help Me! (cos Θ − cos Θ)2 + (cos Θ + cos Θ)2 sin2 Θ 4cos2 Θ 8 cos2 Θ

OpenStudy (anonymous):

(cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (mathmale):

Hello! Please, share the instructions that came with this problem. Also, plese explain what you mean by those '2's. (OK: I see you've realized tht you must use '^' to denote exponentiation.)

OpenStudy (mathmale):

What is your goal in this problem? what have you thought of doing first?

OpenStudy (anonymous):

thats the question and choices

OpenStudy (anonymous):

simplfy both (cos Θ − cos Θ)^2

OpenStudy (mathmale):

surely there are instructions that came with the question.

OpenStudy (mathmale):

Yes. You want to 'simplify' the given expression. What have you considered doing first?

OpenStudy (anonymous):

turning it into cos^2 theta -cos^2 theta to both side

OpenStudy (anonymous):

would it equal sin^2 theta + sin^2 theta

OpenStudy (mathmale):

Your (cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2 doesn't have two sides; what you're trying to say is that this expression HAS TWO TERMS. Let me answer your question with a question: What is (a-b)^2? Think before you answer this. What is (a+b)^2?

OpenStudy (anonymous):

a^2-ab-ba+b^2

OpenStudy (anonymous):

a^2+ab+ba+b^2

OpenStudy (anonymous):

@xapproachesinfinity is the the answer d

OpenStudy (xapproachesinfinity):

It is b, what (a-a)^2 and (a+a)^2?

OpenStudy (mathmale):

Please simplify: a^2-ab-ba+b^2 a^2+ab+ba+b^2 these are called "special products" and you should know these formulas. Once you've simplified and have both formulas correct, apply them to (cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (anonymous):

how did you get that?

OpenStudy (mathmale):

Please do not deal in answers alone. In the end it is up to @helpmeprettyplease4 to find his/her own answers, with our guidance.

OpenStudy (xapproachesinfinity):

(cos Θ − cos Θ)^2 is the same as saying (a-a)^2

OpenStudy (mathmale):

Also, "what (a-a)^2 and (a+a)^2?" is an inappropriate question, mathematically incorrect.l

OpenStudy (mathmale):

Regarding "(cos Θ − cos Θ)^2 is the same as saying (a-a)^2": I disagree.

OpenStudy (mathmale):

@HelpMePrettyPlease4 : Please simplify: a^2-ab-ba+b^2 a^2+ab+ba+b^2 these are called "special products" and you should know these formulas. Once you've simplified and have both formulas correct, apply them to (cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (xapproachesinfinity):

@mathmale my bad! and I agree that the step i did is incorrect to approach this. I just realized it! thanks for the correction. I usually don't give answers actually

OpenStudy (mathmale):

@xapproachesinfinity : Thank you for this mature and honest reply. Since @HelpMePrettyPlease4 has asked for your help, would you mind guiding him/her towards solving his/her original problem?

OpenStudy (anonymous):

a^2-ab+b^2 a^2+ab+b^2

OpenStudy (mathmale):

@xapproachesinfinity : Your input?

OpenStudy (xapproachesinfinity):

it is my pleasure! thanks again

OpenStudy (mathmale):

As explained earlier to @HelpMePrettyPlease4 : I need to get off OpenStudy. I will be available again sometime later today, if needed.

OpenStudy (anonymous):

ok thanks for your help @mathmale :-)

OpenStudy (tkhunny):

I do not understand why any binomial is being considered. cos(x) - cos(x) = 0 -- Square it? Still zero. cos(x) + cos(x) = 2cos(x) -- Square it? 4cos^2(x) Done.

OpenStudy (xapproachesinfinity):

No that is a incorrect as he said! I agree

OpenStudy (anonymous):

ohhh wow thats looks way easier!

OpenStudy (xapproachesinfinity):

that was my initial mistake that i used to approach this

OpenStudy (anonymous):

wait @tkhunny is wrong?

OpenStudy (xapproachesinfinity):

No it easy but incorrect way of doing it. i have done the same thing to get B, and i told you the result right?

OpenStudy (anonymous):

yeah, so what is the correct way to do it?

OpenStudy (tkhunny):

No, tkhunny is not wrong. There is no "right" or "correct" way to do it. Try to simplify your life, when you can. Unique answers don't care how you find them.

OpenStudy (xapproachesinfinity):

so let's go back to the original thing he said that (a-b)^2= a^2-ab-ba+b^2 and (a+b)^2=a^2+ab+ba+b^2 then we simplify (a-b)^2=a^2-2ab+b^2 (a+b)^2=a^2+2ab+b^2 so we use this the solve that

OpenStudy (anonymous):

@tkhunny so the answer is 4cos^2 Θ

OpenStudy (tkhunny):

Like I said, you can do it that way if you like, there is just 500% more to go wrong. Be very, very careful and you can find the right answer that is much more simply generated by a few simple observations. Stop asking for the answer. PROVE it, mark it, and move on. You tell me what the right answer is.

OpenStudy (xapproachesinfinity):

@tkhunny , I have been told it is incorrect! let's discuss that on a post later! with everyone involved

OpenStudy (xapproachesinfinity):

I'm not so sure about it! haha

OpenStudy (anonymous):

ty

OpenStudy (tkhunny):

Awesome. I'm glad you listened when you were told it was wrong. You should have argued. The fact is, cos(x) - cos(x) = 0. There is no adverse consequence to the substitution. Pick an angle: \((\cos(\theta) - cos(\theta))^{2} = 0^{2} = 0\) There is NO counterexample to this conclusion.

OpenStudy (xapproachesinfinity):

you will get cos^2-2cos+cos^2 +cos^2+2cos^2+cos^2 -2cos^2 + 2cos^2 cancels we are left with cos^2+cos^2+cos^2+cos^2

OpenStudy (tkhunny):

Again, if you want to multiply it out, go right ahead. If you are careful, you will reach the same, unique result.

OpenStudy (anonymous):

ty so much! :-) i appreciate all of you guys help!

OpenStudy (xapproachesinfinity):

well I heard the argument, but I don't remember what it was haha.

OpenStudy (aum):

I agree with tkhunny. cos(x) - cos(x) = 0 for all x. But the converse is not always true: cos(A) - cos(B) = 0 does not necessarily imply A = B.

OpenStudy (tkhunny):

And it's not really an argument. No one is talking about A = B in some general sense or any implication about A and B not equal. The problem statement assumes only A = A. There is, therefore, no additional discussion to be had.

OpenStudy (xapproachesinfinity):

You closed it all! @tkhunny, thanks! it is good to have this kind of discussion. I will re check about this and see. I have seen something similar to this. I could be wrong

OpenStudy (aum):

Not everything has to be an argument. The first paragraph of my previous reply applies to this problem. The second paragraph is just an FYI for folks because I have seen the mistaken conclusion made by some.